OpenStudy (anonymous):

What is the slope of the line passing through (0,3) and (4,5)? 0.5 2 -2 -0.5

hero (hero):

(5-3)/(4-0) = 2/4 = 1/2 = 0.5

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1)